# Electric Dipole Moments in 5+3 Flavor Weak Effective Theory

### v1

@@ Appendix A Additional Formulae
$C_{\widetilde{G}}$ $(-7.4\pm 4.5)\cdot 10^{-22}$ $[C_{uG}]_{22}$ $(-15.2\pm 8.2)\cdot 10^{-21}$ $[C_{dG}]_{33}$ $(-10.7\pm 6.2)\cdot 10^{-22}$
$[C_{dG}]_{11}$ $(20.7\pm 9.1)\cdot 10^{-18}$ $[C_{u\gamma}]_{22}$ $(7.3\pm 3.5)\cdot 10^{-22}$ $[C_{d\gamma}]_{33}$ $(-2.0\pm 1.0)\cdot 10^{-23}$
$[C_{dG}]_{22}$ $(15.1\pm 7.6)\cdot 10^{-21}$ $[C^{S1,RR}_{uu}]_{2222}$ $(11.1\pm 5.6)\cdot 10^{-25}$ $[C^{V1,LR}_{uddu}]_{1331}$ $(-12.3\pm 9.5)\cdot 10^{-23}$
$[C_{uG}]_{11}$ $(4.8\pm 1.9)\cdot 10^{-17}$ $[C^{S8,RR}_{uu}]_{2222}$ $(-2.5\pm 1.2)\cdot 10^{-25}$ $[C^{V8,LR}_{uddu}]_{1331}$ $(-34.2\pm 6.0)\cdot 10^{-23}$
$[C_{d\gamma}]_{11}$ $(23.4\pm 1.7)\cdot 10^{-18}$ $[C^{S1,RR}_{uu}]_{1122}$ $(5.6\pm 2.9)\cdot 10^{-22}$ $[C^{V1,LR}_{dd}]_{1331}$ $(1.1\pm 3.2)\cdot 10^{-23}$
$[C_{d\gamma}]_{22}$ $(3.0\pm 2.2)\cdot 10^{-19}$ $[C^{S8,RR}_{uu}]_{1122}$ $(3.7\pm 8.4)\cdot 10^{-23}$ $[C^{V8,LR}_{dd}]_{1331}$ $(17.1\pm 3.0)\cdot 10^{-23}$
$[C_{u\gamma}]_{11}$ $(-89.2\pm 3.5)\cdot 10^{-18}$ $[C^{S1,RR}_{ud}]_{2211}$ $(2.8\pm 1.3)\cdot 10^{-22}$ $[C^{V1,LR}_{dd}]_{2332}$ $(10.7\pm 6.7)\cdot 10^{-25}$
$[C^{V1,LR}_{uddu}]_{1111}$ $(8.7\pm 4.5)\cdot 10^{-22}$ $[C^{S8,RR}_{ud}]_{2211}$ $(12.0\pm 3.7)\cdot 10^{-23}$ $[C^{V8,LR}_{dd}]_{2332}$ $(14.3\pm 8.9)\cdot 10^{-25}$
$[C^{V8,LR}_{uddu}]_{1111}$ $(13.9\pm 6.8)\cdot 10^{-22}$ $[C^{S1,RR}_{ud}]_{2222}$ $(1.0\pm 1.5)\cdot 10^{-26}$ $[C^{S1,RR}_{dd}]_{3333}$ $(2.0\pm 1.0)\cdot 10^{-25}$
$[C^{V1,LR}_{uddu}]_{1221}$ $(16.3\pm 7.6)\cdot 10^{-22}$ $[C^{S8,RR}_{ud}]_{2222}$ $(5.1\pm 3.2)\cdot 10^{-25}$ $[C^{S8,RR}_{dd}]_{3333}$ $(-4.2\pm 2.0)\cdot 10^{-26}$
$[C^{V8,LR}_{uddu}]_{1221}$ $(2.6\pm 1.4)\cdot 10^{-21}$ $[C^{S1,RR}_{uddu}]_{2112}$ $(-13.3\pm 5.5)\cdot 10^{-22}$ $[C^{S1,RR}_{dd}]_{1133}$ $(3.7\pm 1.9)\cdot 10^{-22}$
$[C^{V1,LR}_{dd}]_{1221}$ $(7.5\pm 3.6)\cdot 10^{-22}$ $[C^{S8,RR}_{uddu}]_{2112}$ $(-3.7\pm 5.5)\cdot 10^{-23}$ $[C^{S8,RR}_{dd}]_{1133}$ $(9.5\pm 6.5)\cdot 10^{-23}$
$[C^{V8,LR}_{dd}]_{1221}$ $(12.1\pm 6.7)\cdot 10^{-22}$ $[C^{S1,RR}_{uddu}]_{2222}$ $(-2.4\pm 1.3)\cdot 10^{-24}$ $[C^{S1,RR}_{dd}]_{2233}$ $(12.3\pm 7.0)\cdot 10^{-26}$
$[C^{S1,RR}_{uu}]_{1111}$ $(-13.3\pm 6.5)\cdot 10^{-22}$ $[C^{S8,RR}_{uddu}]_{2222}$ $(-1.9\pm 1.1)\cdot 10^{-24}$ $[C^{S8,RR}_{dd}]_{2233}$ $(-3.7\pm 2.2)\cdot 10^{-25}$
$[C^{S8,RR}_{uu}]_{1111}$ $(4.1\pm 2.2)\cdot 10^{-22}$ $[C^{S1,RR}_{uu}]_{1221}$ $(-1.8\pm 1.3)\cdot 10^{-21}$ $[C^{S1,RR}_{dd}]_{1331}$ $(-2.3\pm 1.1)\cdot 10^{-21}$
$[C^{S1,RR}_{dd}]_{1111}$ $(-6.1\pm 3.4)\cdot 10^{-22}$ $[C^{S8,RR}_{uu}]_{1221}$ $(82.6\pm 9.7)\cdot 10^{-23}$ $[C^{S8,RR}_{dd}]_{1331}$ $(5.2\pm 1.5)\cdot 10^{-22}$
$[C^{S8,RR}_{dd}]_{1111}$ $(1.9\pm 1.0)\cdot 10^{-22}$ $[C^{V1,LR}_{uddu}]_{2112}$ $(-33.2\pm 8.4)\cdot 10^{-24}$ $[C^{S1,RR}_{dd}]_{2332}$ $(14.7\pm 8.5)\cdot 10^{-25}$
$[C^{S1,RR}_{dd}]_{2222}$ $(10.2\pm 5.2)\cdot 10^{-26}$ $[C^{V8,LR}_{uddu}]_{2112}$ $(-4.5\pm 1.0)\cdot 10^{-23}$ $[C^{S8,RR}_{dd}]_{2332}$ $(2.9\pm 1.9)\cdot 10^{-24}$
$[C^{S8,RR}_{dd}]_{2222}$ $(12.1\pm 6.0)\cdot 10^{-26}$ $[C^{V1,LR}_{uddu}]_{2222}$ $(-2.1\pm 1.1)\cdot 10^{-25}$ $[C^{S1,RR}_{ud}]_{1133}$ $(8.0\pm 3.6)\cdot 10^{-22}$
$[C^{S1,RR}_{dd}]_{1122}$ $(-13.7\pm 7.7)\cdot 10^{-23}$ $[C^{V8,LR}_{uddu}]_{2222}$ $(-8.3\pm 4.4)\cdot 10^{-25}$ $[C^{S8,RR}_{ud}]_{1133}$ $(3.9\pm 1.2)\cdot 10^{-22}$
$[C^{S8,RR}_{dd}]_{1122}$ $(5.0\pm 2.3)\cdot 10^{-23}$ $[C^{V1,LR}_{uu}]_{1221}$ $(12.3\pm 2.1)\cdot 10^{-23}$ $[C^{S1,RR}_{uddu}]_{1331}$ $(-6.2\pm 2.5)\cdot 10^{-21}$
$[C^{S1,RR}_{dd}]_{1221}$ $(-22.1\pm 9.4)\cdot 10^{-23}$ $[C^{V8,LR}_{uu}]_{1221}$ $(26.3\pm 2.1)\cdot 10^{-23}$ $[C^{S8,RR}_{uddu}]_{1331}$ $(-1.2\pm 2.8)\cdot 10^{-22}$
$[C^{S8,RR}_{dd}]_{1221}$ $(6.0\pm 2.6)\cdot 10^{-23}$ $[C^{S1,RR}_{ud}]_{2233}$ $(-5.8\pm 3.1)\cdot 10^{-25}$
$[C^{S1,RR}_{ud}]_{1111}$ $(-4.8\pm 2.5)\cdot 10^{-22}$ $[C^{S8,RR}_{ud}]_{2233}$ $(-17.6\pm 9.4)\cdot 10^{-26}$
$[C^{S8,RR}_{ud}]_{1111}$ $(14.7\pm 8.1)\cdot 10^{-23}$ $[C^{S1,RR}_{uddu}]_{2332}$ $(2.8\pm 1.5)\cdot 10^{-24}$
$[C^{S1,RR}_{ud}]_{1122}$ $(-3.0\pm 1.8)\cdot 10^{-22}$ $[C^{S8,RR}_{uddu}]_{2332}$ $(-3.0\pm 1.4)\cdot 10^{-25}$
$[C^{S8,RR}_{ud}]_{1122}$ $(11.4\pm 5.2)\cdot 10^{-23}$ $[C^{V1,LR}_{uddu}]_{2332}$ $(7.5\pm 3.4)\cdot 10^{-26}$
$[C^{S1,RR}_{uddu}]_{1111}$ $(-4.9\pm 2.4)\cdot 10^{-22}$ $[C^{V8,LR}_{uddu}]_{2332}$ $(-4.9\pm 2.4)\cdot 10^{-26}$
$[C^{S8,RR}_{uddu}]_{1111}$ $(14.6\pm 7.5)\cdot 10^{-23}$
$[C^{S1,RR}_{uddu}]_{1221}$ $(-5.0\pm 1.9)\cdot 10^{-22}$
$[C^{S8,RR}_{uddu}]_{1221}$ $(9.7\pm 5.2)\cdot 10^{-23}$
@@
Table 16: Master Formula for $d_{p}$ in WET in the JMS basis at the EW scale. The $\alpha_{I}^{p}$ have units of e-cm.$\rm TeV^{2}$ and e-cm.$\rm TeV$ for the 4-Fermi and dipole operators, respectively.

### v2

@@ Appendix A Additional Formulae
$C_{\widetilde{G}}$ $(-7.6\pm 3.8)\cdot 10^{-22}$ $[C_{uG}]_{22}$ $(-1.1\pm 0.5)\cdot 10^{-20}$ $[C_{dG}]_{33}$ $(-1.5\pm 0.8)\cdot 10^{-21}$
$[C_{dG}]_{11}$ $(2.1\pm 1.0)\cdot 10^{-17}$ $[C_{u\gamma}]_{22}$ $(1.0\pm 0.5)\cdot 10^{-21}$ $[C_{d\gamma}]_{33}$ $(-2.9\pm 1.4)\cdot 10^{-23}$
$[C_{dG}]_{22}$ $(1.5\pm 0.9)\cdot 10^{-20}$ $[C^{S1,RR}_{uu}]_{2222}$ $(7.8\pm 3.9)\cdot 10^{-25}$ $[C^{V1,LR}_{uddu}]_{1331}$ $(-1.5\pm 0.7)\cdot 10^{-21}$
$[C_{uG}]_{11}$ $(4.8\pm 1.8)\cdot 10^{-17}$ $[C^{S8,RR}_{uu}]_{2222}$ $(-1.8\pm 0.9)\cdot 10^{-25}$ $[C^{V8,LR}_{uddu}]_{1331}$ $(-8.1\pm 2.0)\cdot 10^{-22}$
$[C_{d\gamma}]_{11}$ $(2.3\pm 0.2)\cdot 10^{-17}$ $[C^{S1,RR}_{uu}]_{1122}$ $(5.7\pm 2.9)\cdot 10^{-22}$ $[C^{V1,LR}_{dd}]_{1331}$ $(-6.1\pm 3.7)\cdot 10^{-22}$
$[C_{d\gamma}]_{22}$ $(3.0\pm 1.8)\cdot 10^{-19}$ $[C^{S8,RR}_{uu}]_{1122}$ $(3.8\pm 8.2)\cdot 10^{-23}$ $[C^{V8,LR}_{dd}]_{1331}$ $(5.1\pm 1.0)\cdot 10^{-22}$
$[C_{u\gamma}]_{11}$ $(-9.3\pm 0.4)\cdot 10^{-17}$ $[C^{S1,RR}_{ud}]_{2211}$ $(2.9\pm 1.3)\cdot 10^{-22}$ $[C^{V1,LR}_{dd}]_{2332}$ $(1.1\pm 0.6)\cdot 10^{-24}$
$[C^{V1,LR}_{uddu}]_{1111}$ $(5.9\pm 2.9)\cdot 10^{-22}$ $[C^{S8,RR}_{ud}]_{2211}$ $(1.2\pm 0.4)\cdot 10^{-22}$ $[C^{V8,LR}_{dd}]_{2332}$ $(4.1\pm 2.2)\cdot 10^{-24}$
$[C^{V8,LR}_{uddu}]_{1111}$ $(8.9\pm 4.7)\cdot 10^{-22}$ $[C^{S1,RR}_{ud}]_{2222}$ $(1.5\pm 1.4)\cdot 10^{-26}$ $[C^{S1,RR}_{dd}]_{3333}$ $(3.0\pm 1.5)\cdot 10^{-25}$
$[C^{V1,LR}_{uddu}]_{1221}$ $(1.1\pm 0.5)\cdot 10^{-21}$ $[C^{S8,RR}_{ud}]_{2222}$ $(5.1\pm 2.8)\cdot 10^{-25}$ $[C^{S8,RR}_{dd}]_{3333}$ $(-6.4\pm 3.1)\cdot 10^{-26}$
$[C^{V8,LR}_{uddu}]_{1221}$ $(1.7\pm 0.8)\cdot 10^{-21}$ $[C^{S1,RR}_{uddu}]_{2112}$ $(-1.3\pm 0.5)\cdot 10^{-21}$ $[C^{S1,RR}_{dd}]_{1133}$ $(3.8\pm 1.9)\cdot 10^{-22}$
$[C^{V1,LR}_{dd}]_{1221}$ $(5.1\pm 2.7)\cdot 10^{-22}$ $[C^{S8,RR}_{uddu}]_{2112}$ $(-3.6\pm 5.8)\cdot 10^{-23}$ $[C^{S8,RR}_{dd}]_{1133}$ $(9.6\pm 5.3)\cdot 10^{-23}$
$[C^{V8,LR}_{dd}]_{1221}$ $(7.8\pm 4.0)\cdot 10^{-22}$ $[C^{S1,RR}_{uddu}]_{2222}$ $(-2.4\pm 1.5)\cdot 10^{-24}$ $[C^{S1,RR}_{dd}]_{2233}$ $(1.3\pm 0.8)\cdot 10^{-25}$
$[C^{S1,RR}_{uu}]_{1111}$ $(-1.3\pm 0.7)\cdot 10^{-21}$ $[C^{S8,RR}_{uddu}]_{2222}$ $(-1.9\pm 1.2)\cdot 10^{-24}$ $[C^{S8,RR}_{dd}]_{2233}$ $(-3.7\pm 2.4)\cdot 10^{-25}$
$[C^{S8,RR}_{uu}]_{1111}$ $(4.1\pm 2.6)\cdot 10^{-22}$ $[C^{S1,RR}_{uu}]_{1221}$ $(-1.8\pm 1.2)\cdot 10^{-21}$ $[C^{S1,RR}_{dd}]_{1331}$ $(-2.3\pm 1.1)\cdot 10^{-21}$
$[C^{S1,RR}_{dd}]_{1111}$ $(-6.2\pm 3.6)\cdot 10^{-22}$ $[C^{S8,RR}_{uu}]_{1221}$ $(8.3\pm 1.3)\cdot 10^{-22}$ $[C^{S8,RR}_{dd}]_{1331}$ $(5.2\pm 1.6)\cdot 10^{-22}$
$[C^{S8,RR}_{dd}]_{1111}$ $(1.9\pm 0.9)\cdot 10^{-22}$ $[C^{V1,LR}_{uddu}]_{2112}$ $(-2.5\pm 0.9)\cdot 10^{-22}$ $[C^{S1,RR}_{dd}]_{2332}$ $(1.5\pm 0.9)\cdot 10^{-24}$
$[C^{S1,RR}_{dd}]_{2222}$ $(1.0\pm 0.6)\cdot 10^{-25}$ $[C^{V8,LR}_{uddu}]_{2112}$ $(-1.5\pm 0.3)\cdot 10^{-22}$ $[C^{S8,RR}_{dd}]_{2332}$ $(2.9\pm 1.7)\cdot 10^{-24}$
$[C^{S8,RR}_{dd}]_{2222}$ $(1.2\pm 0.6)\cdot 10^{-25}$ $[C^{V1,LR}_{uddu}]_{2222}$ $(-8.1\pm 4.9)\cdot 10^{-25}$ $[C^{S1,RR}_{ud}]_{1133}$ $(8.2\pm 3.7)\cdot 10^{-22}$
$[C^{S1,RR}_{dd}]_{1122}$ $(-1.4\pm 0.9)\cdot 10^{-22}$ $[C^{V8,LR}_{uddu}]_{2222}$ $(-2.8\pm 1.6)\cdot 10^{-24}$ $[C^{S8,RR}_{ud}]_{1133}$ $(3.9\pm 1.1)\cdot 10^{-22}$
$[C^{S8,RR}_{dd}]_{1122}$ $(5.1\pm 2.3)\cdot 10^{-23}$ $[C^{V1,LR}_{uu}]_{1221}$ $(-2.9\pm 18.4)\cdot 10^{-23}$ $[C^{S1,RR}_{uddu}]_{1331}$ $(-6.2\pm 2.5)\cdot 10^{-21}$
$[C^{S1,RR}_{dd}]_{1221}$ $(-2.2\pm 0.9)\cdot 10^{-22}$ $[C^{V8,LR}_{uu}]_{1221}$ $(9.4\pm 0.6)\cdot 10^{-22}$ $[C^{S8,RR}_{uddu}]_{1331}$ $(-1.1\pm 2.8)\cdot 10^{-22}$
$[C^{S8,RR}_{dd}]_{1221}$ $(6.0\pm 2.6)\cdot 10^{-23}$ $[C^{S1,RR}_{ud}]_{2233}$ $(-2.1\pm 1.2)\cdot 10^{-25}$
$[C^{S1,RR}_{ud}]_{1111}$ $(-4.8\pm 2.5)\cdot 10^{-22}$ $[C^{S8,RR}_{ud}]_{2233}$ $(-6.8\pm 3.3)\cdot 10^{-26}$
$[C^{S8,RR}_{ud}]_{1111}$ $(1.5\pm 0.8)\cdot 10^{-22}$ $[C^{S1,RR}_{uddu}]_{2332}$ $(1.4\pm 0.6)\cdot 10^{-24}$
$[C^{S1,RR}_{ud}]_{1122}$ $(-3.0\pm 1.7)\cdot 10^{-22}$ $[C^{S8,RR}_{uddu}]_{2332}$ $(-1.7\pm 0.9)\cdot 10^{-25}$
$[C^{S8,RR}_{ud}]_{1122}$ $(1.1\pm 0.5)\cdot 10^{-22}$ $[C^{V1,LR}_{uddu}]_{2332}$ $(3.8\pm 1.8)\cdot 10^{-25}$
$[C^{S1,RR}_{uddu}]_{1111}$ $(-4.9\pm 2.5)\cdot 10^{-22}$ $[C^{V8,LR}_{uddu}]_{2332}$ $(-1.0\pm 0.5)\cdot 10^{-25}$
$[C^{S8,RR}_{uddu}]_{1111}$ $(1.5\pm 0.8)\cdot 10^{-22}$
$[C^{S1,RR}_{uddu}]_{1221}$ $(-5.1\pm 1.8)\cdot 10^{-22}$
$[C^{S8,RR}_{uddu}]_{1221}$ $(9.7\pm 5.3)\cdot 10^{-23}$
@@
Table 16: Master Formula for $d_{p}$ in WET in the JMS basis at the EW scale. The $\alpha_{I}^{p}$ have units of e-cm.$\rm TeV^{2}$ and e-cm.$\rm TeV$ for the four-fermion and dipole operators, respectively.